A hot air balloonist, rising vertically with a constant velocity of magnitude 5 m/s, releases a sandbag at an instant when the balloon is 40 m above the ground. after it release, the sandbag is in free fall.
a) Complete the position and velocity of the sandbag at 0.250s and 1.00s after its release .
b) How many seconds after its release will the bag strike the ground ?
c) What is the magnitude of velocity does it sytrike the ground? (d) what is the greatest height above the ground that the sandbags reaches (ay-t), vy-t and y-t graph for the motion
1
Expert's answer
2020-03-19T11:36:50-0400

"a=g=9.8 m/s \\
Y_0=40m\\ 
V_0=5m/s"

a) "V=V_0-at"

"Y = Y_0+V_0t-at^2\/2"

velocity at "t=0.250s:"

"V=5-9,8*0.250=2.55 \\ V(t=0.250s)=2.55 m/s"

position at "t=0.250s:"

"Y = 40+5*0.250-\frac{9.8}{2}*0.25^2\\
 Y(t=0.250s)=40.94 m"

velocity at "t=1.00s:"

"V=5-9,8*1.00=4.8 \\ V(t=1.00s)=-4.8 m/s"

position at "t=1.00s:"

Y(t=1.00s)=40.1 m

"Y = 40+5*1.00\frac{9.8}{2}*1.00^2\\
 Y(t=1.00s)=40.1 m"

b)  bag strike the ground at Y=0

"Y =40+5t-4.9t^2\\
40+5t-4.9t^2=0\\
t_1=3.41 \\"

"t_2=-2.3921" - the time can't be below zero

"t=3.41s"


c)

"V=5-9,8*3.41=28.44
 \\ V(t=3.41s)=28.44 m/s"


d) "Y_{max}" if "V=0"

"V=5-9.8t\\
5-9.8t=0\\
t=0.51\\
Y=40+5*0,51-4.9*0.51^2=41.27\\
Y_{max}=41.27 m"